### v1

@@ Abstract
Ensemble Kalman filter with multiple data assimilations; Rotation based iterative Gaussianisation; Geostatistics; Senseor observations; Reconciliation
@@ 1 Introduction
Real-time production data from various sensors can quickly update resource knowledge and inform short-term mine planning decisions. Sensor data provides indirect measurements, known as soft data, which can vary in certainty. Incorporating these observations into geostatistical models can enhance the accuracy and precision of forecasts. Buxton and Benndorf (2013) suggest that integrating sensor data could reduce uncertainty and deviations from production targets, potentially resulting in an average economic benefit of \$5 million per year for the studied deposit.
@@
One practical approach for performing multivariate rapid updating is to transform co-regionalised variables into independent factors before data assimilation. This transformation helps preserve multivariate relationships by back-transforming realisations after updates. Techniques such as the minimum-maximum autocorrelation factors (MAF) (Desbarats and Dimitrakopoulos, 2000) and flow transformation (FT) (van den Boogaart et al, 2017) have been used alongside the EnKF to decorrelate multivariate data before rapid updating (Kumar et al, 2020; Prior et al, 2021b). However, MAF is not well-suited for handling complex multivariate relationships, and FT tends to be too slow for adequate rapid updating. In a comparison of various multivariate transformations, Abulkhair et al (2023) found that the projection pursuit multivariate transform (PPMT) (Barnett et al, 2014) and rotation-based iterative Gaussianisation (RBIG) (Laparra et al, 2011) are considerably faster and more appropriate for rapid updating than FT.
@@
In addition to the EnKF, several other methods have been employed for the rapid updating of resource models, including conditional simulation of successive residuals (Vargas-Guzmán and Dimitrakopoulos, 2002), direct sequential simulation using point distributions (Neves et al, 2019), a variation of the KF for downscaling resource models (Li et al, 2021), and deep reinforcement learning (Kumar and Dimitrakopoulos, 2022). Most of these methods perform single data assimilation, which can lead to some deviations between predicted and observed values, especially when observations are highly uncertain. To improve data assimilation in highly non-linear situations, iterative forms of EnKF have been widely explored in petroleum engineering and hydrogeology. For instance, ensemble randomised maximum likelihood, also known as iterative EnKF (Chen and Oliver, 2012; Gu and Oliver, 2007), yields better-matching results than standard EnKF, although it requires more computational resources. Alternatively, EnKF with multiple data assimilations (EnKF-MDA) (Emerick and Reynolds, 2012, 2013) performs multiple data assimilations with an inflated measurement error to significantly outperform single data assimilation methods. While both iterative EnKF and EnKF-MDA are more computationally intensive than standard EnKF, the EnKF-MDA approach typically requires fewer iterations.
@@
Validating the performance of data assimilation algorithms through real case studies is crucial, especially in the mining industry, where data often show significant variability in terms of spacing between data points, measurement volumes, and associated uncertainties. Several studies of rapid resource model updating have demonstrated the effectiveness of various proposed approaches when applied to real data (Kumar et al, 2020; Prior et al, 2021a, b; Wambeke et al, 2018; Yüksel et al, 2017). In this paper, we apply a combination of EnKF-MDA and RBIG to a real anonymised mining dataset provided by Petra Data Science. The following section offers a detailed methodology for the proposed rapid updating algorithm. Next, real fused data is used to sequentially update the resource model over 25 time periods, the results of which are thoroughly analysed. The paper then concludes with a discussion of key results, limitations of the proposed algorithm, and future research directions.
@@ 2.1 Rapid updating of multivariate resource models
The proposed updating approach uses EnKF-MDA (Emerick and Reynolds, 2012, 2013) for data assimilation, paired with RBIG (Laparra et al, 2011) for decorrelation and Gaussianisation of cross-correlated variables. The steps in the proposed algorithm are as follows:
@@
    $G_{e}=\Phi_{\text{RBIG}}\left(V_{e}\right),$ (1)
    where $V_{e}$ is a vector combining $n$ neighbourhood realisations $Z_{e}^{n,t}$ and observations $O_{e}^{t}$ with the number of variables $e$ at a time $t$, $G_{e}$ is a transformed vector later divided into multi-Gaussian realisations ${Z_{e}^{n,t}}^{\text{RBIG}}$ and observations ${O_{e}^{t}}^{\text{RBIG}}$.
@@
    ${Z_{e}^{n,t+1}}^{\text{RBIG}}=\Phi_{\text{EnKF-MDA}}\left({Z_{e}^{n,t}}^{\text{RBIG}},{O_{e}^{t}}^{\text{RBIG}}\right).$ (2)
@@
    $Z_{e}^{n,t+1}=\Phi_{\text{RBIG}}^{-1}\left({Z_{e}^{n,t+1}}^{\text{RBIG}}\right),$ (3)
    where $\Phi_{\text{RBIG}}^{-1}$ is an inverse RBIG transformation and $Z_{e}^{n,t+1}$ are back-transformed updated neighbourhood realisations.
@@ 2.2 Rotation based iterative Gaussianisation
RBIG is an iterative multi-Gaussian transform based on marginal Gaussianisation followed by orthonormal PCA rotation (Laparra et al, 2011). RBIG is chosen as the optimal multi-Gaussian transform for rapid updating because it is significantly faster than other methods while maintaining comparable performance across various categories (Abulkhair et al, 2023). In a similar study, RBIG demonstrated an approximate 90% reduction in run time compared to PPMT (Cook et al, 2023).
A single iteration of RBIG is defined as
@@
Marginal Gaussianisation functions and rotation matrices at each iteration are stored as $\Phi_{\text{RBIG}}$. The back-transformation $\Phi_{\text{RBIG}}^{-1}$ is performed by following the stored iterations in reverse order.
@@ 2.3 Ensemble Kalman filter with multiple data assimilations
EnKF-MDA is an iterative version of EnKF as it performs multiple data assimilations on the same data with an inflated measurement error (Emerick and Reynolds, 2012). It provided a significantly better match between predictions and observations compared to single data assimilation methods while not being overly computationally intensive (Emerick and Reynolds, 2013). The steps of the EnKF-MDA implemented for rapid resource model updating are the following:
@@
    Repeat steps 1-4 for a predefined number of data assimilations and ensure that random noise added to observations differs from previous data assimilations.
@@ 3.1 Overview of a case study
In this case study, the proposed approach is applied to real data provided by Petra Data Science. The dataset, which has been anonymised, was fused using the company’s Maxta software to track the collected data points back to their respective resource blocks. For more details about data fusion and orebody learning in the MAXTA software, readers can refer to the paper by Petra Data Science (Stewart et al, 2022). Due to confidentiality reasons, we cannot disclose the name of the deposit, the variable names, or the coordinates. Figure 2 provides a 3D view of the study area, specifically highlighting a prior resource model and observations that will be used to update the model. There are nine cross-correlated assay variables, and 18,295 observations are categorised into 25 periods based on the months in which they were collected. For the sake of simplicity, the measurement error is assumed to be a constant value of 10%.
Figure 2: 3D view of a prior resource model with average values of Assay 1 (left) and observations divided into 25 periods (right).
@@ 3.2 An illustrative example of rapid updating for period 1
Table 1: MSE reduction (%) for different numbers of data assimilations in period 1
@@
Predictions versus observations plots indicate that the updated multi-Gaussian results closely align with the diagonal line (Figure 5). After back-transformation, all variables, except for Assay 2 and Assay 8, maintained an error reduction close to 90%. The lower error reductions for these two variables can be attributed to the skewness of their distributions, which impacted the back-transformation process. Notably, although Assay 3 is also highly skewed, its prior prediction is much closer to the diagonal line compared to Assay 2 and Assay 8. The long tails of the predicted values for Assay 2 and Assay 8 contain numerous highly inaccurate predictions, ultimately impacting the final updated results.
@@ 3.3 Sequential rapid updating for the remaining periods
To evaluate the effectiveness of the proposed rapid updating algorithm over time, the resource model realisations were updated sequentially using observations from 25 time periods. The following figures illustrate how the model is progressively refined as more observations become available. Instead of presenting the updates for each individual period, the updates are grouped into approximately equal batches of observations, focusing on a 2D section of the model at an elevation of 44 m.
@@
A more detailed analysis of the final updated models is presented in Figure 13, where predictions are plotted against the observed values. The updated results closely align with the diagonal line, and most variables achieve an error reduction between 86% and 97%. However, Assay 8 only reduced its MSE by 76.73%, which is significantly lower compared to the other variables. This discrepancy is particularly evident in the plot, with many points positioned far from the diagonal line.
A similar trend was noted in Figure 5, where, for period 1, Assay 8 achieved an error reduction of less than 70%. As mentioned earlier, this issue can be partly attributed to the skewness of the distribution, which makes it challenging to accurately back-transform the tail end. Another possible reason for this underperformance is the relatively high number of prior predictions that were significantly inaccurate.
@@
Interestingly, Assay 2 showed an improvement in error reduction compared to period 1, but there are still points that deviate from the diagonal line. Assay 3 demonstrates even better accuracy despite having a more skewed distribution. The difference between Assay 8 and Assay 3 is that the latter has fewer observations at the tail end of its distribution. However, RBIG still struggled to back-transform the tails in both cases, raising concerns about its reliability for highly skewed distributions. Overall, the proposed algorithm effectively reduced the MSE by 94-98% across five out of nine variables.

### v2

@@ Abstract
Ensemble Kalman filter with multiple data assimilations; Rotation based iterative Gaussianisation; Geostatistics; Sensor observations; Reconciliation
@@ 1 Introduction
Real-time production data from various sensors can quickly update resource knowledge and inform short-term mine planning decisions. Incorporating these observations into geostatistical models can enhance the accuracy and precision of forecasts. Buxton and Benndorf (2013) suggest that integrating sensor data could reduce uncertainty and deviations from production targets, potentially resulting in an average economic benefit of \$5 million per year for the studied deposit.
@@
One practical approach for performing multivariate rapid updating is to transform co-regionalised variables into independent factors before data assimilation. This transformation helps preserve multivariate relationships by back-transforming realisations after updates. Techniques such as the minimum-maximum autocorrelation factors (MAF) (Desbarats and Dimitrakopoulos, 2000) and flow transformation (FA) (van den Boogaart et al, 2017) have been used alongside the EnKF to decorrelate multivariate data before rapid updating (Kumar et al, 2020; Prior et al, 2021b). However, MAF is not well-suited for handling complex multivariate relationships, and FA tends to be too slow for adequate rapid updating. In a comparison of various multivariate transformations, Abulkhair et al (2023) found that the projection pursuit multivariate transform (PPMT) (Barnett et al, 2014) and rotation-based iterative Gaussianisation (RBIG) (Laparra et al, 2011) are considerably faster and more appropriate for rapid updating than FA. However, it should be noted that only FA can successfully minimise artefacts in the presence of extreme values, making it a more optimal multi-Gaussian transform, provided that runtime is not important (Cook et al, 2023).
RBIG is similar in its methodology to PPMT, but instead of iteratively searching for interesting projections, it applies orthonormal rotations. As it needs fewer iterations to achieve convergence and each iteration is faster, RBIG is, on average, 90% faster than PPMT when applied to five variables with the number of samples between 10,000 and 50,000 (Cook et al, 2023). Similar to PPMT and FA, RBIG can be paired with other transforms, such as MAF (Erten and Deutsch, 2021; Abulkhair et al, 2023), log-ratio transforms (van den Boogaart et al, 2017; Prior et al, 2021b), and non-logarithmic ratio transforms (Adeli et al, 2017; Bassani et al, 2018).
@@
In addition to the EnKF, several other methods have been employed for the rapid updating of resource models, including conditional simulation of successive residuals (Vargas-Guzmán and Dimitrakopoulos, 2002), direct sequential simulation using point distributions (Neves et al, 2019), a variation of the KF for downscaling resource models (Li et al, 2021), and deep reinforcement learning (Kumar and Dimitrakopoulos, 2022). Most of these methods perform single data assimilation, which can lead to some deviations between predicted and observed values, especially when observations are highly uncertain. To improve data assimilation in highly non-linear situations, iterative forms of EnKF have been widely explored in petroleum engineering and hydrogeology. For instance, ensemble randomised maximum likelihood, also known as iterative EnKF (Gu and Oliver, 2007; Chen and Oliver, 2012), yields better-matching results than standard EnKF, although it requires more computational resources. Alternatively, EnKF with multiple data assimilations (EnKF-MDA) (Emerick and Reynolds, 2012, 2013) performs multiple data assimilations with an inflated measurement error to significantly outperform single data assimilation methods. While both iterative EnKF and EnKF-MDA are more computationally intensive than standard EnKF, the EnKF-MDA approach typically requires fewer iterations. The effectiveness of EnKF-MDA has been demonstrated in hydrogeology (Li et al, 2018), petroleum engineering (Emerick and Reynolds, 2012, 2013) and geophysics (Liu et al, 2023). However, to our knowledge, EnKF-MDA has yet to be applied to the rapid resource model updating problem in mining.
Validating the performance of data assimilation algorithms through real case studies is crucial, especially in the mining industry, where data often show significant variability in terms of spacing between data points, measurement volumes, and associated uncertainties. Several studies of rapid resource model updating have demonstrated the effectiveness of various proposed approaches when applied to real data (Yüksel et al, 2017; Wambeke et al, 2018; Kumar et al, 2020; Prior et al, 2021a, b). However, some of these studies sample observations from a ground truth model generated from real data (Prior et al, 2021a, b). This makes it difficult to account for preferential sampling and incoming new observations from previously under-sampled locations. Both of these problems complicate multivariate transformations since using the same transformation function for model realisations and new observations may become unreliable.
In this paper, we apply a combination of EnKF-MDA and RBIG to a real case study from an iron ore mine in Western Australia. Because of its iterative nature, EnKF-MDA achieves a better match between observations and model-based predictions than EnKF. Additionally, RBIG enables EnKF-MDA to account for multivariate relationships between cross-correlated variables. Because the real data were fused to track historical data to the resource model locations, it makes it easier to validate the proposed approach.
@@
The following section offers a detailed methodology for the proposed multivariate rapid updating algorithm. Next, real fused data is used to sequentially update the resource model over 25 time periods, the results of which are thoroughly analysed. The paper then concludes with a discussion of key results, limitations of the proposed algorithm, and future research directions.
@@ 2.1 Rapid updating of multivariate resource models
The proposed updating approach uses EnKF-MDA for data assimilation, paired with RBIG for decorrelation and Gaussianisation of cross-correlated variables. The steps in the proposed algorithm are as follows:
@@
    $\left({Z_{e}^{n,t}}^{\text{RBIG}},{O_{e}^{t}}^{\text{RBIG}}\right)=\Phi_{\text{RBIG}}^{t}\left(Z_{e}^{n,t},O_{e}^{t}\right),$ (1)
    where $Z_{e}^{n,t}$ are $n$ neighbourhood realisations, $O_{e}^{t}$ are observations with a number of variables $e$ at a time $t$, and $\left({Z_{e}^{n,t}}^{\text{RBIG}},{O_{e}^{t}}^{\text{RBIG}}\right)$ are multi-Gaussian realisations and observations.
@@
    ${Z_{e}^{n,t+1}}^{\text{RBIG}}=\Phi_{\text{EnKF-MDA}}^{t}\left({Z_{e}^{n,t}}^{\text{RBIG}},{O_{e}^{t}}^{\text{RBIG}}\right).$ (2)
@@
    $Z_{e}^{n,t+1}={\Phi_{\text{RBIG}}^{t}}^{-1}\left({Z_{e}^{n,t+1}}^{\text{RBIG}}\right),$ (3)
    where ${\Phi_{\text{RBIG}}^{t}}^{-1}$ is an inverse RBIG transformation and $Z_{e}^{n,t+1}$ are back-transformed updated neighbourhood realisations.
@@ 2.2 Rotation based iterative Gaussianisation
RBIG is an iterative multi-Gaussian transform based on marginal Gaussianisations followed by orthonormal rotations (Laparra et al, 2011). RBIG is chosen as the optimal multi-Gaussian transform for rapid updating because it is significantly faster than other methods while maintaining comparable performance across various metrics (Abulkhair et al, 2023). Principal component analysis (PCA) and independent component analysis (ICA) can be used for orthonormal rotations. Nevertheless, RBIG achieves successful convergence regardless of the choice of orthonormal rotations and although ICA requires fewer iterations, each of its iterations is more computationally complex than those in PCA (Laparra et al, 2011). ICA also generates some significant artefacts after back-transformation, whereas RBIG with PCA produces results similar to PPMT (Abulkhair et al, 2023).
In this paper, PCA is chosen for orthonormal rotations. A single iteration of RBIG is defined as
@@
Marginal Gaussianisation functions and rotation matrices at each iteration are stored as $\Phi_{\text{RBIG}}^{t}$. The back-transformation ${\Phi_{\text{RBIG}}^{t}}^{-1}$ is performed by following the stored iterations in reverse order.
@@ 2.3 Ensemble Kalman filter with multiple data assimilations
EnKF-MDA is an iterative version of EnKF as it performs multiple data assimilations on the same data with an inflated measurement error (Emerick and Reynolds, 2012). It provides a significantly better match between predictions and observations compared to single data assimilation methods while not being overly computationally intensive (Emerick and Reynolds, 2013). The steps of the EnKF-MDA implemented for rapid resource model updating are the following:
@@
    Repeat steps 1-5 for a predefined number of data assimilations and ensure that random noise added to observations differs from previous data assimilations.
@@ 3.1 Overview of a case study
In this case study, the proposed approach is applied to real anonymised mining data. The dataset was fused to track the historical data points back to their respective resource blocks. Readers are referred to the paper (Stewart et al, 2022) for more details about the data fusion and orebody learning involved in creating this dataset. Table 1 shows the statistics of the fused dataset, including the mean, standard deviation, skewness and kurtosis. The data are from an iron ore deposit in Western Australia and include the nine cross-correlated assay variables Fe, S, SiO2, CaO, MgO, P, Al2O3, TiO2 and K2O. Note that the assay variables are not listed in the same order as the variable names in the table. Due to confidentiality reasons, we cannot disclose the name of the deposit, the variable names, or the coordinates.
Table 1: Statistical parameters of the original data.
[…]
The original data points are categorised into 31 periods based on the months in which they were collected. A total of 5,327 data points from the first six months were used to simulate the prior resource model. Turning bands simulation was used to model the multi-Gaussian factors of the original assay variables (the variogram models are shown in Table 2). The remaining 25 periods serve as observations to sequentially update the resource model. Figure 2 provides a 3D view of the study area, specifically highlighting the initial six months of samples, the prior resource model, and the observations that are used to update the model. For simplicity, the measurement error is assumed to be a constant value of 10%.
Table 2: Omni-directional direct variogram models of the multi-Gaussian factors.
[…]
Figure 2: 3D view of initial samples of Assay 1, a prior resource model with average simulated values of Assay 1 and observations divided into 25 periods.
@@ 3.2 An illustrative example of rapid updating for period 1
Table 3: MSE reduction (%) for different numbers of data assimilations in period 1.
@@
Predictions versus observations plots indicate that the updated multi-Gaussian results closely align with the diagonal line (Figure 5). After back-transformation, all variables, except for Assay 2 and Assay 8, maintained an error reduction close to 90%. The lower error reductions for these two variables can be attributed to the combination of different factors. As shown in Table 1 and Figure 5, these two have larger values compared to other variables, are skewed and were predicted less accurately in the prior model. Although Assay 1 also has relatively larger values than others, its distribution is not skewed. On the other hand, despite it being significantly more skewed, Assay 3 has an excellent error reduction of 89% due to a much closer prior prediction. The long tails of the predicted values for Assay 2 and Assay 8 contain numerous highly inaccurate predictions, ultimately impacting the final updated results. Notably, removing Assay 2 and Assay 8 doesn’t significantly affect other variables, with only a 1.24% average difference compared to the results in Figure 5.
@@ 3.3 Sequential rapid updating for the remaining periods
To evaluate the effectiveness of the proposed rapid updating algorithm over time, the resource model realisations were updated sequentially using observations from 25 time periods. However, unlike the rapid updating in period 1, the subsequent periods have one key difference. The neighbourhood around the observations in these periods may contain blocks that were accurately updated in earlier periods. One way to mitigate the issue of potentially overwriting the previously updated blocks is to incorporate observations from prior periods. Therefore, after selecting the neighbourhood around new observations, the proposed algorithm also includes previous observations that are located in the same neighbourhood. Table 4 shows the MSE reduction results after updating the first five periods, comparing the scenarios with and without the inclusion of previous observations. Across all nine variables, the error reductions show consistent improvement. Additionally, since updates are performed within local neighbourhoods around new observations, including previous observations does not significantly slow down the process.
Table 4: MSE reduction (%) after period 5 with and without incorporating previous observations.
[…]
The following figures illustrate how the model is progressively refined as more observations become available. Instead of presenting the updates for each individual period, the updates are grouped into approximately equal batches of observations, focusing on a 2D section of the model at an elevation of 44 m.
@@
A more detailed analysis of the final updated models is presented in Figure 13, where predictions are plotted against the observed values. The updated results closely align with the diagonal line, and most variables achieve an error reduction between 86% and 9%. However, Assay 8 only reduced its MSE by 76.73%, which is significantly lower compared to the other variables. This discrepancy is particularly evident in the plot, with many points positioned far from the diagonal line.
@@
A similar trend was noted in Figure 5, where, for period 1, Assay 8 achieved an error reduction of less than 70%. As mentioned earlier, this issue can be partly attributed to the skewness of the distribution, which makes it challenging to accurately back-transform the tail end. Other factors are the relatively high number of significantly inaccurate prior predictions and the larger value range.
Interestingly, Assay 2 showed an improvement in error reduction compared to period 1, but there are still points that deviate from the diagonal line. Assay 3 demonstrates even better accuracy despite having a more skewed distribution. The difference between Assay 8 and Assay 3 is that the latter has fewer observations at the tail end of its distribution. However, RBIG still struggled to back-transform the tails in both cases, raising concerns about its reliability for highly skewed distributions. Overall, the proposed algorithm effectively reduced the MSE by 94-98% across six out of nine variables.
@@
4 Discussion
The application of the proposed combination of EnKF-MDA and RBIG in a real case study demonstrated its effectiveness in rapidly and accurately updating multivariate resource models. Across all nine variables, the proposed approach achieved an error reduction ranging from 77% to 98%, reducing errors by more than 94% for five of those variables. The results indicate that the updated models provide a more realistic representation of spatial variability and align closely with observations while also maintaining multivariate relationships. Moreover, transforming realisations and observations together helps to account for potential new observations coming from previously under-sampled locations.
In addition to the accuracy, the proposed algorithm operates with impressive speed, even on low-cost hardware. For example, it can update 100 block model realisations with 878 observations in under five minutes. The process could be further enhanced through parallelisation across multiple virtual machines, which is easily achievable in an industrial setting. Notably, the rapid updating requires no human intervention and can be automated to run either at specified time intervals or immediately when new information becomes available. As a result, mining operations can make near real-time decisions based on updated models, leading to improved short-term mine planning and optimisation.
Achieving such computational efficiency and flexibility comes with some drawbacks and limitations. Despite its advantages the Kalman filter (KF) is less effective than kriging or cokriging estimates. Li et al. (Li et al, 2021) compared the updates from KF to estimates produced by ordinary kriging, which unsurprisingly showed better performance for kriging. Incorporating kriging to update resource model realisations can be done by using residuals between observations and prior realisations (Vargas-Guzmán and Dimitrakopoulos, 2002; Chevalier et al, 2015). However, such kriging updates are still computationally demanding and inflexible to be used in the real-time mining framework presented in Figure 1. On the other hand, data assimilation methods are faster, especially when dealing with large resource models. The use of EnKF-MDA helps to further minimise the deviation between model predictions and observations by updating the same data multiple times.
Another drawback noted in this paper is the less accurate updating of extreme values in highly skewed distributions. This limitation is characteristic of RBIG and similar transforms, which can produce artefacts in the presence of extreme values. On the contrary, FA has previously been used for rapid updating (Prior et al, 2021b) and can minimise artefacts after back transformation. However, it is also significantly more computationally intensive and using it for time-sensitive tasks such as rapid updating is not recommended (Abulkhair et al, 2023; Cook et al, 2023). Additionally, multi-Gaussian transforms require the multivariate dataset to be homotopic, making the proposed approach impractical for datasets where some variables are under-sampled. In such cases, a data imputation technique such as Gibbs sampling can be used to generate additional observations (Barnett and Deutsch, 2015; Madani and Bazarbekov, 2021).
Finally, the real data used in the case study has already undergone data fusion and ore tracking. For confidentiality reasons, measurement errors and ore tracking uncertainty were not disclosed. For simplicity, a measurement error of 10% is assumed in this paper. However, the uncertainty associated with sensor observations extends beyond measurement errors typically identified in laboratory settings. The way sensors are used on an actual mine site can differ from how they are intended to be used. Not to mention the difference between harsh mining conditions and a laboratory environment in which measurement errors are evaluated. Furthermore, although more practical, downstream sensors, such as those on conveyor belts, are hard to link back to resource models. This challenge is particularly relevant in underground mining, where ore tracking remains a significant issue. Thus, the uncertainty introduced by ore tracking must be considered during real rapid updating.